INAF queue observing, 2016 July 1 - 2

Observer: F. Cusano, A. Rossi
Telescope Operator: G. Bechetti
Support Astronomer: O Kuhn on call
Instrument: MODS, LUCI

Summary

At the sunset the sky is clear. Some clouds at N-W (UT2:25)

Night duration (12\xB0 el twilight) 20:30-4:00 ( 03:30-11.00 UT) = 7h 30m

Observing time: *0 * (0%)

Weather time loss: 7.5 h (100%)

technical time loss * 0h* (00%)

Data Summary

Project_RUN Objects Filter ExpTime Notes
Standard MODS hz44 dual grating   1 out of 3 expsoures done

Details

04:00 the TO opens the dome

04:20 Humidity at 90%

04:30 slew to standard of Della Ceca, acqBinoMODS Standard_11B.acq mods2r.20160702.0002-0004; mods1r.20160702.0002-0004

04:39 execBinoMODsTarget_11B.obs, mods2r.0005, mods2b.0003; mods1r.0005, mods1b.0003 we have to stop the script, Humidity climbed to 97%, we toke one spectra of the standard, we have to close the dome

07:10 the TO opens the dome humidity 83%

07:25 Humidity at 92%, we try to slew to Testa_A

-- %USERSIG{AndreaRossi - 2016-07-01}%

Comments

 
Topic revision: r5 - 15 Dec 2017, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback